three persons mr.roy, mr.poul and mr.singh together have Rs 51. Mr.poul has Rs 4 less than mr. Roy and mr.singh has got Rs 5 less than mr. Roy . How much money receive by mr.roy,mr.poul and mr.singh.
Answer by Maths68(1474)

Let
Mr. Roy has = x
Mr. Pual has = x-4
Mr. Singh has = x-5
Together they have = 51
Mr. Roy + Mr. Pual + Mr. Singh = 51
x+(x-4)+(x-5)=51
3x-9=51
3x=51+9
3x=60
divide by 3
3x/3=60/3
x=20

Mr. Roy has = x = 20
Mr. Pual has = x-4 = 20-4 = 16
Mr. Singh has = x-5 = 20-5= 15
20+16+15=51
51=51